The human circulation system has approximately 1 ✕ 109 capillary vessels. Each vessel has a diameter of about 9 µm. Assuming cardiac output is 6 L/min, determine the average velocity (in cm/s) of blood flow through each capillary vessel.
Answer #1

Area, A = (pi/4) d^2

A = (pi/4) x (9 x 10^-6)^2 = 6.36 x 10^-11 m^2

Total area, A' = 10^9 x 6.36 x 10^-11 = 0.064 m^2

Velocity, v = (6 x 10^-3)/(0.064 x 60) = 0.156 cm/s

Comment in case any doubt please rate my answer....